The Magellanic Clouds Newsletter

⋆ ⋆ ⋆

An electronic exchange on Magellanic Clouds research

Edited by Bryan C. Dunne Co-Editors Eva K. Grebel and You-Hua Chu

                               mcnews@astro.uiuc.edu
                http://www.astro.uiuc.edu/projects/mcnews/MCNews.php

No. 76 May 25th, 2004

Letter From The Editor

Dear Colleagues,

The Magellanic Clouds Working Group is pleased to bring you issue 76 of the Magellanic Clouds Newsletter. In this issue, we present 4 submitted abstracts, an announcement for an upcoming meeting in honor of Carl Heiles, and a collection of recent astro-ph listings of potential interest to Magellanic Clouds researchers. As always, this month’s issue is available from the MC News Website in a variety of formats (PDF, HTML, PostScript, and L^{A}TEX).

Best Regards, Bryan Dunne Editor, MC News

Contents

3 Abstracts of Refereed Papers . . . . . . . . . . . . . 2 1 Poster Session Abstract . . . . . . . . . . . . . . . . . . . 4 1 Meeting Announcement . . . . . . . . . . . . . . . . . . . 4 Recent astro-ph Listings . . . . . . . . . . . . . . . . . . . . 6

Upcoming Meeting Quicklist

The Nature and Evolution of Disks Around Hot Stars . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . July 7–9

Astrophysics in the Far Ultraviolet: Five Years of Discovery with FUSE . . . . . . . . . August 2–6

Astrophysics and Cosmology After Gamow - Theory and Observations . . . . . . . August 8–14

Diffuse Matter in the Galaxy: Observations Confront Theory . . . . . . . . . . . . . . Aug 29–Sept 2

Abstracts of Refereed Papers

Multi-wavelength analysis of the dust emission in the Small Magellanic Cloud

C. Bot (1 and 2), F. Boulanger (2), G. Lagache (2), L. Cambresy (1), D. Egret (3)

(1) Observatoire astronomique de Strasbourg, Strasbourg, France

(2) Institut d’Astrophysique Spatiale (IAS), Orsay, France

(3) Observatoire de Paris, Paris, France

We present an analysis of dust grain emission in the diffuse interstellar medium of the Small Magellanic Cloud (SMC). This study is motivated by the availability of 170 microns ISOPHOT data covering a large part of the SMC, with a resolution enabling to disentangle the diffuse medium from the star forming regions. After data reduction and subtraction of Galactic foreground emission, we used the ISOPHOT data together with HiRes IRAS data and ATCA/Parkes combined HI column density maps to determine dust properties for the diffuse medium. We found a far infrared emissivity per hydrogen atom 30 times lower than the Solar Neighborhood value. The modeling of the spectral energy distribution of the dust, taking into account the enhanced interstellar radiation field, gives a similar conclusion for the smallest grains (PAHs and very small grains) emitting at shorter wavelength. Assuming Galactic dust composition in the SMC, this result implies a difference in the gas-to-dust ratio (GDR) 3 times larger than the difference in metallicity. This low depletion of heavy elements in dust could be specific of the diffuse ISM and not apply for the whole SMC dust if it results from efficient destruction of dust by supernovae explosions.

Comments: Accepted by A&A WWW: http://arxiv.org/abs/astro-ph/0405045 e-mail: bot@astro.u-strasbg.fr

Discovery of a New Pulsar Wind Nebula in the Large Magellanic Cloud

B. M. Gaensler (1), S. P. Hendrick (2), S. P. Reynolds (1 and 2), K. J. Borkowski (2)

(1) Harvard-Smithsonian Center for Astrophysics

(2) North Carolina State University

We present new high-resolution radio and X-ray observations of the supernova remnant (SNR) B0453-685 in the Large Magellanic Cloud, carried out with the Australia Telescope Compact Array and the Chandra X-ray Observatory respectively. Embedded in the SNR shell is a compact central nebula producing both flat-spectrum polarized radio emission and non-thermal X-rays; we identify this source as a pulsar wind nebula (PWN) powered by an unseen central neutron star. We present a new approach by which the properties of a SNR and PWN can be used to infer upper limits on the initial spin period and surface magnetic field of the unseen pulsar, and conclude that this star was an initial rapid rotator with current properties similar to those of the Vela pulsar. As is the case for other similarly-aged sources, there is currently an interaction taking place between the PWN and the SNR’s reverse shock.

Comments: Published in ApJ, v594, L111 (2003) WWW: http://arxiv.org/abs/astro-ph/0305574 e-mail: bgaensler@cfa.harvard.edu

Quantitative Studies of the Far-UV, UV and Optical Spectra of Late O and Early B-type Supergiants in the Magellanic Clouds

C. J. Evans (1), P. A. Crowther (2), A. W. Fullerton (3,4), D. J. Hillier (5)

(1) Isaac Newton Group of Telescopes, Canary Islands, Spain

(2) Dept. of Physics & Astronomy, University of Sheffield, Sheffield, UK

(3) Dept. of Physics & Astronomy, University of Victoria, Victoria, Canada

(4) Center for Astrophysical Sciences, Dept. of Physics & Astronomy, The Johns Hopkins University, Baltimore, MD, USA

(5) Department of Physics & Astronomy, University of Pittsburgh, Pittsburgh, PA, USA

We present quantitative studies of 8 late O and early B-type supergiants in the Magellanic Clouds using far-ultraviolet F USE, ultraviolet IUE/HST and optical VLT-UVES spectroscopy. Temperatures, mass-loss rates and CNO abundances are obtained using the non-LTE, spherical, line-blanketed model atmosphere code of Hillier & Miller (1998). We support recent results for lower temperatures of OB-type supergiants as a result of stellar winds and blanketing, which amounts to ∼2000 K at B0 Ia. In general, Hα derived mass-loss rates are consistent with UV and far-UV spectroscopy, although from consideration of the S iv λλ1063-1073 doublet, clumped winds are preferred over homogenous models. AV 235 (B0 Iaw) is a notable exception, which has an unusually strong Hα profile that is inconsistent with the other Balmer lines and UV wind diagnostics. We also derive CNO abundances for our sample, revealing substantial nitrogen enrichment, with carbon and oxygen depletion. Our results are supported by comparison with the Galactic supergiant HD 2905 (BC0.7 Ia) for which near-solar CNO abundances are obtained. This bolsters previous suggestions that “normal” OB-type supergiants exhibit atmospheric compositions indicative of partical CNO processing.

Comments: Accepted by ApJ WWW: http://www.ing.iac.es/˜ cje/lateO.ps.gz e-mail: cje@ing.iac.es

Poster Session Abstracts

Structure of the Periphery of the Large Magellanic Cloud Further Revealed by 2MASS

David R. Alves

Columbia University

In order to better understand the nature of equilibrium structures like the flared outer disk and hot stellar halo of the Large Magellanic Cloud (LMC), I have reexamined the content of the 2MASS database on lines of sight toward the periphery of the LMC. I find evidence for a metallicity gradient as a function of projected radius based on the dereddened colors of the LMC’s red giant branch (RGB) stars. I also find that the projected center of the number density contours depends systematically on the selected color (metallicity) of the RGB stars. Implications for microlensing are discussed.

Comments: 8 page PDF (this is the poster, not a conference proceedings) WWW: http://widefield.lbl.gov/2004/posters/alves d.pdf e-mail: alves@astro.columbia.edu

Meeting Announcements

Diffuse Matter in the Galaxy: Observations Confront Theory

August 29–September 2, 2004 Arecibo, Puetro Rico, USA

The interstellar medium (ISM) provides the material to form stars and occupies the vast majority of space. As a result of star formation and stellar energy feedback, the physical conditions of the ISM evolve and produce multiple phases. These include neutral and ionized phases and hot and cold phases. On scales from a few pc to a few hundred pc, interstellar shells form from stellar winds and supernovae. Galaxy-wide, these processes lead to a matrix of ISM having vastly different physical conditions. Moreover, the formation and evolution of all interstellar structures are subject to influences of the magnetic field.

Once the province of optical and radio observations, the ISM is now observed at all wavelengths, from the ground and from space. At the same time, new theoretical techniques are emerging to study the ISM on small (star-forming) scales and on the galactic scale. In this conference, we will examine closely the recent observational and theoretical studies of diffuse ISM in the Galaxy and in others. The goal is to provide an improved understanding of the nature and evolution of this material.

For four decades, Carl Heiles has played a pioneering role in studies of the diffuse ISM in the Galaxy. His work, primarily observational, has strongly influenced theoretical ideas of the medium. Therefore, it is most fitting to honor Carl, a distinguished researcher and an inspiring professor, with a meeting dedicated to observational and theoretical studies of the diffuse ISM.

Scientific Program The scientific program will include observational and theoretical discussions of these topics: ∘ CNM/WNM ∘ Dust ∘ H i/H_{2} relationship ∘ WIM ∘ HIM ∘ The Galactic Disk/Halo interaction ∘ The global nature of the diffuse ISM ∘ Magnetic fields ∘ SNRs, shells & supershells

Our plan is to organize coherent sessions, each session will consider one or more of topics listed above. Each session will be guided and organized in advance by a Chair who will consult with session presenters prior to the meeting. In this way, we hope to achieve a more systematic exposition of recent progress and future directions for each topic considered.

Conference Capacity

We can accommodate up to 100 participants. Early interest in this conference has been high. If you have not yet indicated your interest in attending, please let us know soon. Also, if you indicated your interest in attending but your plans have now changed, please let us know.

Deadlines ∘ Proposed presentation title submission: May 30, 2004 ∘ Early Registration deadline: June 15, 2004

For more information, contact the meeting organizers at: WWW: http://www.naic.edu/meeting.htm E-mail: ism@astro.uiuc.edu

Recent astro-ph Listings

astro-ph/0405412 Title: Metal abundances of RR Lyrae stars in the bar of the Large Magellanic Cloud Authors: R.G. Gratton (1), A. Bragaglia (2), G. Clementini (2), E. Carretta (1,2), L. Di Fabrizio (2,3), M. Maio (2), E. Taribello (4)- ((1) INAF - Oss. Astron. Padova, (2) INAF - Oss. Astron. Bologna, (3) Telescopio Nazionale Galileo, (4) Astron. Dept. Bologna Univ.) Comments: 17 pages, 22 figures, figs 1 to 8 only as JPEG, uses aa.cls, in press on A&A

astro-ph/0405409 Title: New Calibrations of Pulsational Absolute Magnitudes of Field RR Lyrae Stars Using Revised Dependencies of Temperatures, Masses, and Periods on Metallicity Authors: Allan Sandage (Carnegie Observatories, Pasadena, CA) Comments: 20 pages, 4 figures, 3 tables. Submitted to Monthly Notices of the RAS

astro-ph/0405401 Title: A Hubble Space Telescope WFPC2 Investigation of the Disk-Halo Interface in NGC891 Authors: Joern Rossa (STScI), Ralf-Juergen Dettmar (AIRUB), Rene A. M. Walterbos (NMSU), Colin A. Norman (JHU/STScI) Comments: LaTeX, 15 pages, 11 figures, accepted for publication in The Astronomical Journal, paper with high resolution figures can be accessed at this http URL

astro-ph/0405378 Title: OH (1720 MHz) Maser Search Toward the Large Magellanic Cloud Authors: C. L. Brogan, W. M. Goss, J. Lazendic, A. J. Green Comments: 20 pages, 7 figures, Accepted to AJ

astro-ph/0405377 Title: Properties of RR Lyrae stars in the Inner Regions of the Large Magellanic Cloud Authors: J.Borissova, D.Minniti, M.Rejkuba, D.Alves, K.H.Cook, K.C.Freeman Comments: 13 pages, 16 figures, A&A in press

astro-ph/0405299 Title: A near-infrared survey of the N49 region around the Soft Gamma-Ray Repeater 0526-66 Authors: S. Klose, A. A. Henden, U. Geppert, J. Greiner, H. H. Guetter, D. H. Hartmann, C. Kouveliotou, C. B. Luginbuhl, B. Stecklum, F. J. Vrba Comments: submitted to ApJ Letters

astro-ph/0405257 Title: LMC Self-lensing from a new perspective Authors: L. Mancini (1,2,4), S. Calchi Novati (1), Ph. Jetzer (1), G. Scarpetta (2,3,4) ((1) ITP, Zurich;

(2) Dipartimento di Fisica, Universita di Salerno; (3) IIASS, Vietri Sul Mare; (4) INFN, sez. di Napoli)

astro-ph/0405133 Title: A Comprehensive Catalogue of Variable Stars in the field of 47 Tucanae Authors: David T F Weldrake, Penny D Sackett, Terry J Bridges, Kenneth C Freeman Comments: 62 pages, 24 figures. Accepted for publication in AJ. For a full resolution version and variable star findercharts, see: this http URL

astro-ph/0405129 Title: Rotation of Early B-type Stars in the Large Magellanic Cloud - The role of evolution and metallicity Authors: Stefan C. Keller Comments: 15 pages, PASA accepted astro-ph/0405122

Title: Confronting the Superbubble Model with X-ray Observations of 30 Dor C Authors: David A. Smith, Q. Daniel Wang Comments: 25 pages, 4 figures. To appear in The Astrophysical Journal, August 20, 2004 issue

astro-ph/0405045 Title: Multi-wavelength analysis of the dust emission in the Small Magellanic Cloud Authors: C. Bot (1 and 2), F. Boulanger (2), G. Lagache (2), L. Cambresy (1), D. Egret (3) ((1)Observatoire astronomique de Strasbourg, (2)IAS/Orsay, (3)Observatoire de Paris) Comments: 11 pages, 10 figures. Accepted for publication in Astronomy & Astrophysics

astro-ph/0405015 Title: SALT and RR Lyrae variables: Our Galaxy, The Magellanic Clouds and the Local Group Authors: Michael Feast Comments: 10 pages,no figures. Science with SALT workshop, in press

astro-ph/0404561 Title: Long Period Variables in the Magellanic Clouds: OGLE + 2MASS + DENIS Authors: Martin Groenewegen Comments: Astronomy and Astrophysics, accepted

astro-ph/0404547 Title: A Dynamical Model for the Orbit of the Andromeda Galaxy M31 and the Origin of the Local Group of Galaxies Authors: T. Sawa (Department of Physics and Astronomy, Aichi University of Education), M. Fujimoto (Department of Physics and Astrophysics, Nagoya University) Comments: Submitted and revised to A&A, 21 pages, 14 figures and 2 tables

astro-ph/0404523 Title: The Optical Gravitational Lensing Experiment. The New Catalog of Eclipsing Binary Stars in the Small Magellanic Cloud Authors: L. Wyrzykowski, A. Udalski, M. Kubiak, M.K. Szymanski, K. Zebrun, I. Soszynski, P.R. Wozniak, G. Pietrzynski, O. Szewczyk Comments: 15 pages. Figure 3 and Appendixes A-C included in ’jpg’ format. Full resolution paper, Appendices and photometric data presented in the paper are available from the OGLE Internet archive: this http URL or its US mirror: this http URL Journal-ref: Acta Astron. 54 (2004) 1

astro-ph/0404398 Title: Magellanic Cloud Cepheids: Pulsational and Evolutionary Modelling vs. Observations Authors: J. Robert Buchler, Zoltan Kollath, Jean-Philippe Beaulieu Comments: 8 pages, 5 figures. Change of title and minor revisions

astro-ph/0404395 Title: Luminosity Profiles of Resolved Young Massive Clusters Authors: Francois Schweizer (Carnegie Observatories, Pasadena) Comments: 10 pages, LaTeX, 6 figures (= 11 EPS files); uses asp2004.sty. To appear in ”Formation and Evolution of Massive Young Star Clusters,” eds. H.J.G.L.M. Lamers, A. Nota, & L.J. Smith (ASP, San Francisco), 2004

astro-ph/0404376 Title: Evolution of planetary nebulae II. Population effects on the bright cut-off of the PNLF Authors: P. Marigo (1), L. Girardi (2), A. Weiss (3), M.A.T. Groenewegen (4), C. Chiosi (1) ((1) Padova Univ., (2) Trieste Obs., (3) MPA Garching, (4) Leuven Univ.) Comments: 22 pages, to appear in Astronomy & Astrophysics astro-ph/0404367

Title: Central Stars of Planetary Nebulae in the Large Magellanic Cloud: A Far-UV Spectroscopic Analysis Authors: J.E. Herald, L. Bianchi (The Johns Hopkins University) Comments: 53 pages, 15 figures (11 color). Accepted for publication in ApJ

astro-ph/0404362 Title: IR Surface Brightness Fluctuations of Magellanic Star Clusters Authors: Rosa A. Gonzalez (CRyA, UNAM, Morelia, Mexico), Michael C. Liu (IFA, Hawaii, USA), Gustavo Bruzual-A. (CIDA, Merida, Venezuela) Comments: 54 pages, 12 figures, accepted by ApJ

astro-ph/0404300 Title: High mass X-ray binaries in LMC: dependence on the stellar population age and the ”propeller” effect Authors: P. Shtykovskiy (1), M. Gilfanov (2,1) ((1) Space Research Institute, Moscow, Russia; (2) Max-Planck-Institut fuer Astrophysik, Garching, Germany) Comments: 20 pages, 11 figures Submitted to A&A

astro-ph/0404280 Title: A deep survey of heavy element lines in planetary nebulae – II. Recombination line abundances and evidence for ultra-cold plasma Authors: Y. G. Tsamis, M. J. Barlow, X.-W. Liu, P. J. Storey, I. J. Danziger Comments: 27 pages, 7 figures, submitted to the MNRAS

astro-ph/0404250 Title: Candidates for giant lobes projecting from the LBV stars P Cygni and R 143 Authors: J. Meaburn (1), P. Boumis (2), M. P. Redman (3), J. A. Lopez (4), F. Mavromatakis (5) ((1) Jodrell Bank Observatory, University of Manchester, UK, (2) Institute of Astronomy & Astrophysics, National Observatory of Athens, Greece, (3) Dublin Institute for Advanced Studies, School of Cosmic Physics, Ireland, (4) Instituto de Astronomia, UNAM, Mexico, (5) University of Crete, Heraklion, Crete, Greece) Comments: 7 pages, 3 figures, accepted for publication in A&A

astro-ph/0404247 Title: Photometry of Magellanic Cloud clusters with the Advanced Camera for Surveys - I. The old LMC clusters NGC 1928, 1939 and Reticulum Authors: A.D. Mackey, G.F. Gilmore (IoA, Cambridge, UK) Comments: 15 pages, 9 figures. Accepted for publication in MNRAS. For a version with full resolution figures visit this http URL (recommended)